OpenStudy (anonymous):

can anyone help? 3/x=1/x-2 Anwsers: x = 3 x = 4 x = 6

jimthompson5910 (jim_thompson5910):

Cross multiply to get 3(x-2) = 1*x Then solve 3x-6 = x -6 = x - 3x -6 = -2x -6/(-2) = x 3 = x x = 3

jimthompson5910 (jim_thompson5910):

btw I answered that other question in the other post